OpenStudy (anonymous):

What is the distance between points (-3, -7) and (-7, -9)?

OpenStudy (anonymous):

d=squareroot 20

OpenStudy (anonymous):

\[distance=\sqrt{20}\]

OpenStudy (anonymous):

i gotcha!

OpenStudy (anonymous):

\[d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\]

OpenStudy (anonymous):

The distance formula is \(d=\sqrt{(x_{1} - x_{2})^{2} + (y_{1} - y_{2})^{2}}\) \( d = \sqrt{(-3+7)^{2} + (-7+9)^{2}}\) \(d=\sqrt{(4)^{2} + (2)^{2}}\) \(d=\sqrt{16 + 4}\) \(d=\sqrt{20}\)

OpenStudy (anonymous):

너무 단순해서